# The process of polarisation as a loss of dimensionality: measuring changes in polarisation using Singular Value Decomposition of Random Dot Product Graphs

### v3

@@ 2 Methods
2.1 Network modelling
Having established these networks, we model them as Random Dot Product Graphs (RDPGs) [39]. RDPGs are used instead of other graph embeddings because their optimal embedding dimension (see below) is established a priori to the analysis, so it is independent of the network’s size.
In the most general, directed, case under the RDPG model, each node $i\in\{1,...,N\}$ in a graph $G$ is associated with two vectors of traits, $L_{i}$ and $R_{i}$, that give the node position in a pair of metric spaces, $L$ and $R$. $L_{i}$ and $R_{i}$ are in general not directly observable. Then, the probability that an edge from node $i$ to node $j$ exists is given by the proximity of $L_{i}$ and $R_{j}$, namely by the dot product
$L_{i}\cdot R_{j}=\mathbb{P}(i\rightarrow j)\,.$
In other words, the position of a node in $L$ describe its outgoing edge topology, and the position of a node in $R$ describe its incoming edge topology. In general, given the two matrices $L$ and $R$, the edges are drawn with independent probabilities given by $LR$. We call the couple $(L,R)$ the RDPG embedding of $G$.
In inference tasks, starting from an observed graph, the goal is to estimate the position of the nodes in the latent spaces, given the interaction structure of the network. We do not parametrise the network for this analysis. For a fixed dimension $d$ of the two latent spaces, this is achieved by a $d$-truncated Singular Value Decomposition as follows (full description of SVD can be found in Noble [40].
Let $A$ be the adjacency matrix of $G$. Let $A=U\Sigma V^{\prime}$ be a singular value decomposition of A, so that $U$ and $V^{\prime}$ are orthogonal matrices, and $\Sigma$ is the diagonal matrix whose $i$-th entry is the $i$-th singular value of $A$ (sorted in decreasing order). Notice that in general $U$ and $V^{\prime}$ are only identifiable up to orthogonal transformations (any rotation of them would keep the dot product constant, so they would determine the same graph). Denoting $M|_{k}$ the truncation of a matrix $M$ to its first $k$ columns, for any $d$, the two matrices $\hat{L}=U|_{d}\sqrt{\Sigma}$ and $\hat{R}=\sqrt{\Sigma}\left(V|_{d}\right)^{\prime}$ determine a rank-$d$ optimal approximations of $A$. That is, $\hat{L}\hat{R}=\hat{A}$ minimizes the Frobenius distance to $A$ between all the rank-$d$ matrices. ${\Sigma}$ is truncated to the first $d$ diagonal elements; $\hat{L}$ has $N$ rows and $d$ columns, and $\hat{R}$ has $d$ rows and $N$ columns.
@@ 2.1 Network modelling
We define the dimension of a communication network as the optimal choice of $d$ for the RDPG embedding of the network and denote it $\hat{d}$.
An a-priori optimal choice for $\hat{d}$ can be obtained from, $\Sigma$, the sorted sequence of singular values of the network’s adjacency matrix $A$. Various methods exist. Here we adopt the elbow method presented in [41]. The elbow method identifies the most likely change point in the sequence of values of $\Sigma$ by sequentially fitting two Gaussian distributions with independent mean, and equal variance. One Gaussian distribution is fitted to the largest $d$ singular values, and the other to the smallest $K-d$. Then, the optimal $\hat{d}$ is the value of $d$ that maximise the sum of the log-likelihoods of the two distributions.
Notice that $\hat{d}$ is robust to network size, i.e. its value is determined by the complexity of the network rather than the size of the network [39].
@@
Given a network, we can assess its graph complexity by computing its SVD entropy, which choose this instead of other entropy measures because it is based on $\Sigma$. A network has higher SVD entropy when many of its singular vectors are highly important for its structure, meaning that the network cannot be efficiently compressed. This is commonly read as an indication of high network complexity [42], and it is related (although not in a linear nor straightforward way) to its dimension. We normalise the SVD entropy using Pielou’s evenness [43], so that the results do not depend on the network size.
@@
2.2 Polarisation
We define a process of polarisation as the loss of dimensionality of a graph observed in time. Namely, we find the optimal RDPG embedding dimension $\hat{d}$ at multiple time points. If $\hat{d}$ decreases over time then we argue that the network has become more polarised during that time. This is based on the same principles as the view that the process of polarisation is one of increasing correlation. The dimensions of $\hat{d}$ are all uncorrelated; as such reduction in $\hat{d}$ corresponds to a reduction in uncorrelation in the social network graph, which is an increase in correlation.
We complement this definition by also observing the complexity of the graph, as determined by its SVD entropy, and notice whether it corresponds to an increase or decrease of polarisation.
2.3 Code
All the Social Network analysis discussed above have been performed in Julia[44], in particular using the packages Graphs.jl[45] for network manipulation, PROPACK.jl[46] for computing the (truncated) singular values, and DotProductGraphs.jl for computing the embedding dimension and SVD entropy.
All scripts are available at https://doi.org/10.5281/zenodo.11043841. In Julia all packages can be installed by running Pkg.activate(“.”); Pkg.install().
@@ 3.1 Climate discussion in New Zealand Twitter
Figure 1: Plot comparing $\hat{d}$ of NZ climate change tweets in 2017-2020 to $\hat{d}$ in 2020-2023.
Figure 2: Plot comparing the SVD entropy of NZ climate change tweets in 2017-2020 to the entropy in 2020-2023.
@@ 3.2 COP discussion in Twitter
Falkenberg et al. collected a very large corpus of tweets by querying Twitter’s Academic API v2.0 for tweets mentioning ”COP2x” where x was an integer between 0 and 6 (inclusive) [33]. Details about this data can be found in their original paper. They restricted their search to tweets in English, and covered the COP from 20 to 26 (years 2014 to 2022, with 2020 and 2021 skipped due to the Covid-19 pandemic). Their adjacency matrix was constructed based on whether a user $i$ retweeted tweets from a political influencer $j$. Their focus was whether there was noticeable division among users based on whether they were spreading true information about climate change or disinformation from climate change denialist influencers. To test for polarisation using our framework, we built a network for each year’s COP using the same data by considering each user (identified by their unique IDs) as a node, and any mention, reply, or quote tweet between two users as an edge. The network for each year was analysed independently. We then graphed $\hat{d}$ and the SVD entropy for each year in order to show how they, and therefore the polarisation of networks that discussed the COP conferences, changed over time.
[h]
@@
20 54 0.97711855
21 47 0.9803695
22 62 0.9787037
23 52 0.9789476
24 42 0.9741172
25 78 0.9807784
26 38 0.9823967
We found that $\hat{d}$ for the network of Twitter users discussing the COP conference has been decreasing over time, though the decrease was not linear. Unexpectedly, the SVD entropy of the network did not decrease in this way, and instead it was at its highest in 2022 even though $\hat{d}$ was at its lowest.
Interestingly, Falkenberg et al. expected to find polarisation during COP21, due to the signing of the Paris Agreement at COP21 [33]. Their Hartigan’s Dip Test for COP21 returned a significant result (p = 0.003), but they go on to claim that COP21 was not polarised despite this result. In our data, COP21 has a lower $\hat{d}$ than the years before or after. It may be possible that the network as a whole became more polarised, which is captured by our data, but this effect had not yet occurred among the ”influencers” that Falkenberg et al selected. Our results support Falkenberg et al’s suggestion that the increase in polarisation they observed was due to an increase in the prominence of anti-climate and generally far-right influencers on Twitter, since COP26 was the conference with the lowest $\hat{d}$.
@@ 4 Conclusions
We have demonstrated a novel method for measuring polarisation through the embedded dimensionality $\hat{d}$ of random dot-product graphs. This is a reliable and straightforward implementation of the correlation-based approach to polarisation suggested by Baldassarri and Gelman. Our method captured the presence of polarisation in all the scenarios where it was expected and had been found by other researchers, in both simulated data and real social media networks. In particular, the ability to observe the changes in $\hat{d}$ over time in data from the COP conferences provided a useful demonstration of how this method can be applied to longitudinal data. The RDPG approach also allows us to easily see that the process of polarisation is occurring in a network over time, through its embedded dimensionality reducing, rather than relying on a binary test of whether the network was polarised or non-polarised.
Another advantage of the RDPG approach is that it is computationally light; the main bottleneck is the computation of the first singular values of a large matrix, but this is well known in computer science literature and has already been strongly optimised. We found that the SVD was feasible even when used on networks with millions of nodes. Bimodality-based methods typically use SVD or correspondance analysis to determine the dimension they will test for bimodality, so our approach is at least as efficient.
Our approach is also highly interpretable, without forcing the latent ideological distributions into an artificially unidimensional space. Rather than creating a unidimensional space and then interpreting its political meaning (such as pro- and anti-climate, or left- and right-wing), the dimensionality method instead focuses on the number of dimensions rather than what those dimensions are. In high-dimensional spaces, we do not need to know exactly what ideologies the dimensions correspond to; the important part is that they signal that there are ideological connections being made between nodes that would not be possible if the network was polarised.
@@
Another possible extension of this method would be to implement a nonparametric two-sample hypothesis test[47], since this would allow a hypothesis test of whether the two networks are significantly different as additional evidence of polarisation having occurred. We believe that being able to observe the embedded dimensionality of the graph alone is useful; however, we understand that sometimes a hypothesis test is demanded, and we believe this would help demonstrate that changes in the embedded dimensionality of the network are significant. In our tests using stochastic block models, we found an expected result that $\hat{d}$ increases as the probability of edges forming between the groups increases. This is evidence for the common belief about polarisation that connections between the groups decrease their polarisation. We also found an unexpected result that $\hat{d}$ does not straightforwardly increase when one group is much larger than the other, and in fact decreases greatly when one group is 100 times the size of the other. While this is somewhat a property of the size of the network and its sparsity, since the small group was very small and thus unable to make a large number of edges with the big group, we believe that it is an important case to have tested and that closer investigation of this scenario is warranted. We think that properties sparsity are important for modelling the real conditions of social networks at risk of polarisation, such as communities who have recently received refugees, and should be carefully considered in the model design to ensure that real effects are not being mathematically counteracted.
It would also be productive to experiment more with stochastic block models, since we only tested two possible scenarios. For example, this paper has only explored the two-block case; many instances of online ”echo chambers” have a large number of groups who all hate people different to them, and it would be useful to see what happens to the embedded dimensionality in such cases. Similarly, our testing on group prevalence showed a decrease in embedded dimensionality when one group was 100 times the size of the other, but we did not follow the smaller group’s size all the way to zero so we do not know whether the embedded dimensionality remains low or rebounds to a higher value.
4.1 Acknowledgements

### v6

@@ 2.2.2 Inferring an RDPG model
At the time the data was collected in 2021, their terms of service required researchers to only retain the tweed IDs rather than all the data. This was designed to allow researchers to reconstruct the networks for re-analysis as needed without creating data security and copyright problems. However, after Elon Musk’s takeover in 2022, Twitter’s download access was changed to make it extremely expensive – sometimes completely impossible – to download these networks.
@@
We define the dimension of a communication network as the optimal choice of $d$ for the RDPG embedding of the network and denote it $\hat{d}$.An a-priori optimal choice for $\hat{d}$ can be obtained from, $\Sigma$, the sorted sequence of singular values of the network’s adjacency matrix $A$. Various methods exist. Here we adopt the elbow method presented in [46]. The elbow method identifies the most likely change point in the sequence of values of $\Sigma_{k}$ by sequentially fitting two Gaussian distributions with independent mean, and equal variance. One Gaussian distribution is fitted to the largest $d$ singular values, and the other to the smallest $k-d$. Then, the optimal $\hat{d}$ is the value of $d$ that maximise the sum of the log-likelihoods of the two distributions. Notice that $\hat{d}$ is robust to network size, i.e. its value is determined by the complexity of the network rather than the size of the network [44].
@@
Given a network, we can assess its graph complexity by computing its SVD entropy. We choose this instead of other entropy measures because it is based on $\Sigma$, meaning that it will be easier to compare it to the results for the network’s embedded dimensionality than if we had chosen an entropy not based in $\Sigma$. It has also previously been used in ecology to measure the complexity of ecological networks [47], so we think it would be useful to test whether it works for the complexity of social networks. A network has higher SVD entropy when many of its singular vectors are highly important for its structure, meaning that the network cannot be efficiently compressed. This is commonly read as an indication of high network complexity [48], and it is related (although not in a linear nor straightforward way) to its dimension. We normalise the SVD entropy using Pielou’s evenness [49], so that the results do not depend on the network size.
@@
2.3 Data, script and code availability
Our analysis was conducted using the Julia programming language. Scripts, data, and code can be downloaded from https://doi.org/10.5281/zenodo.15180262 and run on a computer that has Julia installed. A README file is included in the repository [50].
@@ 3.1 Climate discussion in New Zealand Twitter
Time Period Tweets Users Edges Giant Component
2017-2020 6767 1402 2258 914
2020-2023 6172 1217 1997 1013
Table 1: Characteristics of NZ Climate Change conversation networks.
@@
Figure 1: Plot comparing $\hat{d}$ of NZ climate change tweets in 2017-2020 to $\hat{d}$ in 2020-2023. $\hat{d}$ has decreased in the second time period, which may indicate that the network is more polarised than in the first time period.
Figure 2: Plot comparing the SVD entropy of NZ climate change tweets in 2017-2020 to the entropy in 2020-2023. The entropy has decreased in the second time period, which may indicate that the network is more polarised than in the first time period
Year Dimension Dimension GC Entropy Entropy GC
2017-2020 39 39 0.980 0.980
2020-2023 27 24 0.974 0.974
@@ 3.2 COP discussion in Twitter
Falkenberg et al. collected a very large corpus of tweets by querying Twitter’s Academic API v2.0 for tweets mentioning ”COP2x” where x was an integer between 0 and 6 (inclusive) [35]. We reconstructed the same networks based on the tweet IDs given in the data released with the paper. Unfortunately due to changes in Twitter’s Acadimic API in 2022, this reconstruction process is now at best prohibitively expensive, and at worst impossible. They restricted their search to tweets in English, and covered the COP from 20 to 26 (years 2014 to 2022, with 2020 and 2021 skipped due to the Covid-19 pandemic). Their adjacency matrix was constructed based on whether a user $i$ retweeted tweets from a political influencer $j$. Their focus was whether there was noticeable division among users based on whether they were spreading true information about climate change or disinformation from climate change denialist influencers. The smallest network was 87,780 users and the largest was 7,809,303. More information about the networks can be found on p.39 of the supplementary materials in [35].To test for polarisation using our framework, we built a network for each year’s COP using the same data by considering each user (identified by their unique IDs) as a node, and any mention, reply, or quote tweet between two users as an edge. Unlike Falkenberg et al. we do not make a distinction between influencers and regular users, in order to capture the effects that interaction between normal users has on the dimensionality of the network. The network for each year was analysed independently. We then graphed $\hat{d}$ and the SVD entropy for each year in order to show how they, and therefore the polarisation of networks that discussed the COP conferences, changed over time.
@@
20 54 0.977
21 47 0.980
22 62 0.979
23 52 0.979
24 42 0.974
25 78 0.981
26 38 0.982
We found that $\hat{d}$ for the network of Twitter users discussing the COP conference has been decreasing over time, though the decrease was not linear. Unexpectedly, the SVD entropy of the network did not decrease in this way, and instead it was at its highest in 2022 even though $\hat{d}$ was at its lowest. To make sure that our results were not being driven by changes in the network size, we tested the Pearson correlation of the $\hat{d}$ values with the number of users in the network. For $\hat{d}$ based on the first 100 singular values the correlation had a p-value of 0.34 and for $\hat{d}$ based on the first 1,000 singular values the correlation had a p-value of 0.3542. We are therefore confident that changes in the size of the networks are not affecting our results. Interestingly, regardless of whether calculation of $\hat{d}$ was restricted to 100 or 1,000 singular values, COP20 had a $\hat{d}$ value the same as or greater than other years despite having a much smaller network, and COP26 had the lowest $\hat{d}$ despite having the largest network. Interestingly, Falkenberg et al. expected to find polarisation during COP21, due to the signing of the Paris Agreement at COP21 [35]. Their Hartigan’s Dip Test for COP21 returned a significant result (p = 0.003), but they go on to claim that COP21 was not polarised despite this result. In our data, COP21 has a lower $\hat{d}$ than the years before or after. It may be possible that the network as a whole became more polarised, which is captured by our data, but this effect had not yet occurred among the ”influencers” that Falkenberg et al selected. Our results support Falkenberg et al’s suggestion that the increase in polarisation they observed was due to an increase in the prominence of anti-climate and generally far-right influencers on Twitter, since COP26 was the conference with the lowest $\hat{d}$.
@@ 4 Conclusions
We have demonstrated a novel method for measuring polarisation using the embedded dimensionality $\hat{d}$ of random dot-product graphs. This is a reliable and straightforward implementation of the correlation-based approach to polarisation suggested by Baldassarri and Gelman [37], and supported with observations from political scientists Laclau and Mouffe [40]. Our method captured the presence of polarisation in all the scenarios where it was expected and had been found by other researchers, in both simulated data and real social media networks. In particular, the ability to observe the changes in $\hat{d}$ over time in data from the COP conferences provided a useful demonstration of how this method can be applied to longitudinal data. The RDPG approach also allows us to easily see that the process of polarisation is occurring in a network over time, through its embedded dimensionality reducing, rather than relying on a binary test of whether the network was polarised or non-polarised.
Another advantage of the RDPG approach is that it is computationally light; the main bottleneck is the computation of the first singular values of a large matrix, but this is well known in computer science literature and has already been strongly optimised. We found that the SVD was feasible even when used on networks with millions of nodes. Bimodality-based methods typically use SVD or correspondence analysis to determine the dimension they will test for bimodality, so our approach is at least as efficient.
Our approach is also highly interpretable, without forcing the latent ideological distributions into an artificially unidimensional space. Rather than creating a unidimensional space and then interpreting its political meaning (such as pro- and anti-climate, or left- and right-wing), the dimensionality method instead focuses on the number of dimensions rather than what those dimensions are. In high-dimensional spaces, we do not need to know exactly what ideologies the dimensions correspond to; the important part is that they signal that there are connections being made between nodes that would not be possible if the network was polarised.
@@
Another possible extension of this method would be to implement a nonparametric two-sample hypothesis test[51], since this would allow a hypothesis test of whether the two networks are significantly different as additional evidence of polarisation having occurred. We believe that being able to observe the embedded dimensionality of the graph alone is useful; however, we understand that sometimes a hypothesis test is demanded, and we believe this would help demonstrate that changes in the embedded dimensionality of the network are significant.
In our tests using stochastic block models, we found an expected result that $\hat{d}$ increases as the probability of edges forming between the groups increases. This is evidence for the common belief about polarisation that connections between the groups decrease their polarisation. We also found an unexpected result that $\hat{d}$ does not straightforwardly increase when one group is much larger than the other, and in fact decreases greatly when one group is 100 times the size of the other. While this is somewhat a property of the size of the network and its sparsity, since the small group was very small and thus unable to make a large number of edges with the big group, we believe that it is an important case to have tested and that closer investigation of this scenario is warranted. We think that allowing the blocks to become sparsely connected without correcting this property is important for modelling the real conditions of social networks at risk of polarisation, such as communities who have recently received a small number of refugees, and models should be carefully designed to ensure that these effects are retained when appropriate.
It would also be productive to experiment more with stochastic block models, since we only tested two possible scenarios. For example, this paper has only explored the two-block case; many instances of online ”echo chambers” have a large number of groups who all hate people different to them, and it would be useful to see what happens to the embedded dimensionality in such cases. Similarly, our testing on group prevalence showed a decrease in embedded dimensionality when one group was 100 times the size of the other, but we did not follow the smaller group’s size all the way to zero so we do not know whether the embedded dimensionality remains low or rebounds to a higher value when one group becomes extremely small.
5 Acknowledgements
A preprint version of this article has been peer-reviewed and recommended by PCI Network Sci (https://doi.org/10.24072/pci.networksci.100217) [52]. We would like to thank them for their excellent process.